The Length of a Rectangle is 2 cm More Than 5 Times the Width. The Perimeter is 196 cm. Find the Width and Length.
Answer by mananth(16946)

The Length of a Rectangle is 2 cm More Than 5 Times the Width. The Perimeter is 196 cm. Find the Width and Length.
...
width = x
length = 5x+2
...
2*(L+W)= perimeter
2*(x+5x+2)= 196
2*(6x+2)=196
12x+4 = 196
add -4 to both sides
12x= 192
/12
x=16 cm. the width
Length = 5x+2 = 82 cm
...
m.ananth@hotmail.ca